what are the dimensions of a rectangle whose perimeter is 22 square meters and its are is 30 square meters?
Answer by richwmiller(17219)

Sorry that is not possible. Perimeter cannot be square meters. Perimeter can be yards, meters , miles , inches

2L+2W=22
l+w=11
l*w=30
l=6 w=5